OpenStudy (anonymous):

\[x \in \mathbb{R}\] Show that \[\frac{2x+1}{6}(2(x+1)(2x+3)-2x(2x-1))=(2x+1)^{2}\]

OpenStudy (mathmath333):

(2(x+1)(2x+3)-2x(2x-1)) =4x^2+10x+6-4x^2+2 =12x+6 =6(2x+1)

OpenStudy (anonymous):

thank you very much i got it

OpenStudy (mathmath333):

y wlcme